a different car accelerated from 0.12m/s to 0.52m/s

The acceleration of the car was 0.040m/s^2

The work done to accelerate the car was 0.48j

calculate the resultant force needed to accelerate the car.


1
Expert's answer
2021-04-26T17:11:40-0400
"v_f^2-v_i^2=2as=2a\frac{W}{F}\\0.52^2-0.12^2=2(0.04)\frac{0.48}{F}\\F=0.15\ N"